OpenStudy (anonymous):

a line AB passe through the point c(-2,3) and is parallel to the line y-1 equal-1/2(x+1). write the equation of line AB in slope-point form

OpenStudy (anonymous):

Because the line AB is parallel to the line : \[y-1=\frac12(x+1)\] then, the two lines have the same slope m which is 1/2. So : \[m=\frac12\] Now, the line AB passe through the point C(-2,3). So we can say : \[y-3=m(x-(-2))\] So : \[y-3=\frac12(x+2)\]

OpenStudy (anonymous):

thank u